### 07 May 2026 v2

# Some geometric series for Euler’s constant

Abstract.
We provide representations of Euler’s constant $\gamma=0.577$... as series which converge geometrically fast (but use coefficients whose computation induces a quadratic cost). The asymptotic oscillations of these coefficients are discussed.
Key words and phrases:
Euler’s constant, Euler-Mascheroni constant, asymptotic of sequences
2020 Mathematics Subject Classification:
11Y60, 11B83 (Primary) 11B68, 11M41, 40G99 (Secondary)
1. Main result
Let $e_{0}=0$, and define positive rational numbers $e_{m}$ for $m\geq 1$ by this recurrence:
$e_{m}=\frac{2^{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}e_{m-j}}{2^{m+1}-2}.$ (1)
This gives the sequence $2$, $\frac{7}{3}$, $\frac{8}{3}$, $\frac{133}{45}$, $\frac{16}{5}$, … , $e_{10}=\frac{163287}{40579}$, …. See Table 1 for more values.
Table 1. The coefficients $e_{m}$, $1\leq m\leq 20$
$m$ $e_{m}$
11 22
22 $\frac{7}{3}$
33 $\frac{8}{3}$
44 $\frac{133}{45}$
55 $\frac{16}{5}$
66 $\frac{3221}{945}$
77 $\frac{3392}{945}$
88 $\frac{100391}{26775}$
99 $\frac{20848}{5355}$
$10$ $\frac{163287}{40579}$
$m$ $e_{m}$
$11$ $\frac{7567072}{1826055}$
$12$ $\frac{10605587147}{2492565075}$
$13$ $\frac{1551804656}{356080725}$
$14$ $\frac{1732332761353}{388911367845}$
$15$ $\frac{252492267136}{55558766835}$
$16$ $\frac{2313623814645529}{499751107680825}$
$17$ $\frac{261522788700176}{55527900853425}$
$18$ $\frac{69661896931499841923}{14556250513419389775}$
1919 $\frac{2828470111061381408}{582250020536775591}$
2020 $\frac{23101294621895391907711}{4689192129680103420375}$
As usual, $\gamma$ is Euler constant. Here is the main Theorem.
Theorem 1.
Let $\ell\geq 2$. Let $(e_{m})$ be the sequence defined by (1). There holds
$\gamma=\sum_{n=1}^{2^{\ell-1}-1}\frac{1}{n}-(\ell-1)\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$ (2)
In particular for $\ell=2$:
$\gamma=1-\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\Bigl(2^{-m-1}+3^{-m-1}\Bigr),$
and for $\ell=3$:
$\gamma=1+\frac{1}{2}+\frac{1}{3}-2\log 2+\sum_{m=1}^{\infty}\frac{(-1)^{m-1}e_{m}}{m+1}\Bigl(4^{-m-1}+5^{-m-1}+6^{-m-1}+7^{-m-1}\Bigr).$
The level $\ell$ is the number of binary digits used by the integers whose inverse powers are added in Equation (2). In Table 2, we illustrate the numerical values obtained using up to $e_{10}$ for $\ell$ being equal to $2$, 33, …, $7$ (so that the computations could have been done by a diligent human calculator), and also up to $e_{20}$ for $\ell=2,3,4$. The computed approximations to Euler’s constant $\gamma$ in this table are truncated not rounded.
Table 2. Table of approximations of $\gamma$
$\ell$ Series from Thm. 1 up to $e_{m}=e_{10}$ last term
2 0.577 15 $\approx\numprint{-0.0001807}$
[…]
$\gamma$ 0.577 215 664 901 532 860 606 512 090
$\ell$ Series from Thm. 1 up to $e_{m}=e_{20}$ last term
[…]
A numerical implementation using Python and the mpmath library is provided at https://burnolmath.gitlab.io/dyadic-gamma/ (it uses by default $\ell=8$). One can express (see Equation (13)) $e_{m}$ as an explicit finite linear combination of the Bernoulli numbers $B_{k}$, $1\leq k\leq m$, but this uses terms which are much larger than $e_{m}$ itself. So, it seems that using the recurrence (1), which is stable numerically, is more appropriate for computing $e_{m}$’s with hundreds of decimal places. However, an algorithm using (1) appears to have an intrinsic (at least) quadratic cost, and is not suited for large scale computations.
[…]
Table 3. Comparison of $e_{m}$ with $(\log 2)^{-1}H_{m+1}$
$m$ $e_{m}-(\log 2)^{-1}H_{m+1}$
11 -0.164 042 561 333
22 -0.311 607 574 963
33 -0.338 948 001 852
44 -0.338 598 121 140
55 -0.334 602 850 177
66 -0.332 236 533 267
77 -0.331 621 032 426
88 -0.331 908 031 990
99 -0.332 424 034 678
$10$ -0.332 833 436 510
$m$ $e_{m}-(\log 2)^{-1}H_{m+1}$
$11$ -0.333 041 375 341
$12$ -0.333 074 427 538
$13$ -0.332 999 205 051
$14$ -0.332 880 028 058
$15$ -0.332 762 836 214
$16$ -0.332 673 218 057
$17$ -0.332 620 451 671
$18$ -0.332 603 053 818
1919 -0.332 613 748 636
2020 -0.332 643 144 549
Figure 1. $e_{m}-(\log 2)^{-1}H_{m+1}$ vs $\log_{2}(m)$, $16\leq m\leq 500$
Theorem 2.
There holds
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m})$ (3)
with
$\phi(t)=-(\log 2)^{-1}\sum_{n\in\mathbb{Z},n\neq 0}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr)\mathsf{e}^{2\pi int}.$
[…]
Figure 2. $e_{m}-(\log 2)^{-1}H_{m+1}$ as function of $\log_{2}(m)$, and the graph of $\frac{\gamma}{\log 2}+\frac{1}{2}+\phi(t)$, $16\leq m\leq 1000$
As a corollary to the Theorem, $e_{2m}\approx 1+e_{m}$, a fact which does not seem immediately obvious from Equation 1.
Remark 1.
Numerical investigations suggest that the logarithm in the numerator of the error term from Equation (3) can be omitted. However we should remain cautious because we can compute in reasonable time only of the order of ten thousand coefficients. To be more precise, the error term appears to have the shape $\psi(\log_{2}(m))/m+o(1/m)$ with some $1$-periodic function $\psi$ averaging to $\frac{3}{2\log(2)}$.
Remark 2.
[…]
Remark 3.
We will establish concrete explicit upper and lower bounds for $e_{m}$. They are useful, especially the upper bound, to calibrate numerical implementations.
2. Geometric series for $\gamma$
We obtain Theorem 1 as a corollary to our previous study of the Euler alternating series [5].
Proof.
In [5, Eq. (5)], a sequence of meromorphic functions $(c_{m}(s))$ , $m\geq 0$, $s\in\mathbb{C}$, is defined by the initial condition $c_{0}(s)=1$ and, for $m\geq 1$:
$c_{m}(s)=\frac{1}{2^{m+s}-2}\sum_{j=1}^{m}\binom{m}{j}c_{m-j}(s).$ (4)
In [5], there is an integer parameter $b>1$, which here has been set to $b=2$. There are generalizations of Theorem 1 to general $b>2$, where the analog of the recurrence from Equation (1) has additional factors given by the power sums of the base-$b$ digits.
Let $\ell\geq 2$, the level, be some integer greater than one. The Euler alternating series $\eta(s)=\sum_{n=1}^{\infty}(-1)^{n-1}n^{-s}$ was proven in [5, Thm. 1] to verify for $\Re s>0$ the next formula, which converges locally uniformly:
$\eta(s)=\frac{2^{s}-2}{2^{s}}\sum_{0<n<2^{\ell-1}}\frac{1}{n^{s}}+\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{s}}+\sum_{m=1}^{\infty}(-1)^{m}\frac{(s)_{m}}{m!}c_{m}(s)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{s+m}}\,.$ (5)
We now compute the derivative at $s=1$ (here $H_{0}=0$):
$\eta^{\prime}(1)=(\log 2)H_{2^{\ell-1}-1}+\sum_{m=0}^{\infty}(-1)^{m}\left\{\begin{aligned} &H_{m}c_{m}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\\ &+c_{m}^{\prime}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\\ &-c_{m}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{\log n}{n^{m+1}}\end{aligned}\right$. (6)
Recall from [5] that $c_{m}(1)=(m+1)^{-1}$ (which is easily checked on the recurrence (4)). Leaving temporarily aside the consideration of the $c_{m}^{\prime}(1)$, we compute:
$\sum_{m=0}^{\infty}(-1)^{m}\biggl(\frac{-1}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{\log n}{n^{m+1}}+\frac{H_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}\biggr)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=1}\sum_{m=1}^{\infty}\frac{(-1)^{m}}{m+1}\frac{(s)_{m}}{m!}(1/n)^{m+1}\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\left.\frac{\mathrm{d}}{\mathrm{d}s}\right|_{s=1}\int_{0}^{\frac{1}{n}}(1+x)^{-s}\mathrm{d}x\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})+\int_{0}^{\frac{1}{n}}\frac{-\log(1+x)}{(1+x)}\mathrm{d}x\right)$
$=\sum_{2^{\ell-1}\leq n<2^{\ell}}\left(-\log n\log(1+\frac{1}{n})-\frac{1}{2}\log^{2}(1+\frac{1}{n})\right)$
$=-\frac{1}{2}\sum_{2^{\ell-1}\leq n<2^{\ell}}\log(1+\frac{1}{n})\log\bigl(n^{2}(1+\frac{1}{n})\bigr)=-\frac{1}{2}\sum_{2^{\ell-1}\leq n<2^{\ell}}\bigl(\log^{2}(n+1)-\log^{2}(n)\bigr)$
$=-\frac{1}{2}\Bigl(\ell^{2}\log^{2}2-(\ell-1)^{2}\log^{2}2\Bigr)=-(\ell-\frac{1}{2})\log^{2}2.$
Hence, from Equation (6):
$\eta^{\prime}(1)=(\log 2)H_{2^{\ell-1}-1}-(\ell-\frac{1}{2})\log^{2}2+\sum_{m=0}^{\infty}(-1)^{m}c_{m}^{\prime}(1)\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$
Now, as is well-known, $\eta^{\prime}(1)=\log(2)(\gamma-\frac{1}{2}\log 2)$, so, with the definition
$e_{m}=-(m+1)\frac{c_{m}^{\prime}(1)}{\log 2},$ (7)
(in particular $e_{0}=0$), we obtain
$\gamma=H_{2^{\ell-1}-1}-(\ell-1)\log 2+\sum_{m=1}^{\infty}(-1)^{m-1}\frac{e_{m}}{m+1}\sum_{2^{\ell-1}\leq n<2^{\ell}}\frac{1}{n^{m+1}}.$
The proof of Theorem 1 will be complete once it has been confirmed that the $(e_{m})$ sequence defined by Equation (7) verifies the recurrence (1). Taking the derivative at $s=1$ of Equation (4), and recalling $c_{m}(1)=(m+1)^{-1}$, we obtain, for $m\geq 1$:
$c_{m}^{\prime}(1)$ $=\frac{-(\log 2)2^{m+1}}{2^{m+1}-2}c_{m}(1)+(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m}{j}c_{m-j}^{\prime}(1)$
$e_{m}$ $=\frac{2^{m+1}}{2^{m+1}-2}+(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m}{j}\frac{m+1}{m-j+1}e_{m-j}.$
This, indeed, is Equation (1). And $e_{0}=0$. ∎
3. Estimates for the coefficients $e_{m}$
In this section, we establish the next proposition as a corollary to the recurrence from Equation (1).
Proposition 1.
For $m\geq 2$, there holds
$\frac{H_{m+1}}{\log 2}-\numprint{0.35}<e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.31}\;.$
We first establish a less tight bound. We will then re-use the tools from the proof of the Lemma 1 to prove Proposition 1.
Lemma 1.
There holds for any $m\in\mathbb{N}$ (with $H_{n}=\sum_{j=1}^{n}j^{-1}$):
$\frac{H_{m+1}-1}{\log 2}\leq e_{m}<\frac{H_{m+1}}{\log 2}-\numprint{0.161}\,.$
Proof.
Let us define $\delta_{n}=e_{n}-(\log 2)^{-1}H_{n+1}$ for $n\geq 0$. We thus have, for $m\geq 1$:
$(2^{m+1}-2)e_{m}=\left\{\begin{aligned} 2^{m+1}&+(\log 2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}\\ &+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.\end{aligned}\right$. (8)
Now (as is well-known):
$H_{m+1}+\sum_{j=1}^{m}\binom{m+1}{j}H_{m-j+1}$ $=\sum_{k=1}^{m+1}\binom{m+1}{k}H_{k}$
$=\int_{0}^{1}\sum_{k=0}^{m+1}\binom{m+1}{k}\frac{1-t^{k}}{1-t}\mathrm{d}t$
$=\int_{0}^{1}\frac{2^{m+1}-(1+t)^{m+1}}{1-t}\mathrm{d}t$
$=\int_{1}^{2}\frac{2^{m+1}-u^{m+1}}{2-u}\mathrm{d}u$
$=2^{m+1}\int_{\frac{1}{2}}^{1}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\int_{0}^{\frac{1}{2}}\frac{1-x^{m+1}}{1-x}\mathrm{d}x$
$=2^{m+1}H_{m+1}-2^{m+1}\log 2+2^{m+1}\underbrace{\int_{0}^{\frac{1}{2}}\frac{x^{m+1}}{1-x}\mathrm{d}x}_{\epsilon_{m}}.$
Hence, from (8),
$(2^{m+1}-2)e_{m}=(2^{m+1}-2)\frac{H_{m+1}}{\log 2}+\frac{H_{m+1}}{\log 2}+2^{m+1}\frac{\epsilon_{m}}{\log 2}+\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}.$
$e_{m}=\frac{H_{m+1}}{\log 2}+\underbrace{\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\frac{\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}}{2^{m+1}-2}}_{\delta_{m}}$ (9)
Suppose now by induction that $\delta_{j}\geq-(\log 2)^{-1}$, for $0\leq j<m$, which is true for $m=1$, as $\delta_{0}=-(\log 2)^{-1}$. The above implies $\delta_{m}\geq-(\log 2)^{-1}$ too, so it is true for all $m$.
To obtain an upper bound, we define
$\mu_{m}=\max_{0\leq j\leq m}\delta_{j}.$
In particular $\mu_{0}=\delta_{0}=-(\log 2)^{-1}\approx\numprint{-1.443}$, and $\mu_{1}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164}$. The above implies, for $m\geq 1$:
$\mu_{m}\leq\frac{H_{m+1}+2^{m+1}\epsilon_{m}}{(2^{m+1}-2)\log 2}+\mu_{m-1},$
hence by induction, we obtain, for any given $m$:
$p>m$ $\implies\mu_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{p}\frac{H_{n+1}+2^{n+1}\epsilon_{n}}{2^{n+1}-2},$
thus, using
$\frac{1}{2^{n+1}-2}$ $=\frac{1}{2^{n+1}}+\frac{2}{2^{n+1}(2^{n+1}-2)}\leq\frac{1}{2^{n+1}}+\frac{1}{4^{n}},$
we obtain, for any $p>m\geq 0$:
$\delta_{p}\leq\mu_{m}+(\log 2)^{-1}\sum_{n=m+1}^{\infty}\left(2^{-n-1}H_{n+1}+4^{-n}H_{n+1}+\epsilon_{n}+2^{-n+1}\epsilon_{n}\right).$
Now,
$\sum_{n=m+1}^{\infty}2^{-n-1}H_{n+1}=\frac{-\log(1-\frac{1}{2})}{1-\frac{1}{2}}-\sum_{n=0}^{m}2^{-n-1}H_{n+1},$
and
$\sum_{n=m+1}^{\infty}4^{-n}H_{n+1}=4\frac{-\log(1-\frac{1}{4})}{1-\frac{1}{4}}-\sum_{n=0}^{m}4^{-n}H_{n+1}.$
Evaluating numerically for $m=10$ we obtain respectively:
$(\log 2)^{-1}\sum_{n=11}^{\infty}2^{-n-1}H_{n+1}$ $\approx\numprint{0.002237}$
[…]
We bound trivially $\epsilon_{n}=\int_{0}^{\frac{1}{2}}x^{n+1}(1-x)^{-1}\mathrm{d}x<2^{-n-1}$, hence:
$(\log 2)^{-1}\sum_{n=11}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-10})\sum_{n=11}^{\infty}2^{-n-1}$
$=(\log 2)^{-1}(2^{-11}+2^{-21})\approx\numprint{0.000705}$
So for $p>10$, $\delta_{p}<\mu_{10}+\numprint{0.003}$. A quick look at Table 3 shows that $\mu_{10}=\delta_{1}=2-(\log 2)^{-1}\frac{3}{2}\approx\numprint{-0.164043}$. Hence $\mu_{10}+\numprint{0.003}<\numprint{-0.161}$. This concludes the proof of Lemma 1. ∎
Proof of Proposition 1.
We continue with the same technique as in the previous proof. To improve the upper bound, we separate out $\delta_{1}$:
$(2^{m+1}-2)^{-1}\sum_{j=1}^{m}\binom{m+1}{j}\delta_{m-j}$ $\leq\frac{m(m+1)}{2(2^{m+1}-2)}\delta_{1}$
$\mkern-100.0mu+\frac{2^{m+1}-2-m(m+1)/2}{2^{m+1}-2}\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
As we know that $\delta_{j}\geq\delta_{0}=-(\log 2)^{-1}$, we can further bound this as
$\leq\frac{m(m+1)}{2(2^{m+1}-2)}\bigl(\underbrace{\delta_{1}+|\delta_{0}|}_{2-\frac{1}{2}(\log 2)^{-1}\approx\numprint{1.279}}\bigr)+\max\{\delta_{j},0\leq j<m,j\neq 1\}.$
We again use $(2^{n+1}-2)^{-1}<2^{-n-1}(1+2^{-n+1})$ for $n\geq 1$. So:
$\sum_{n=21}^{\infty}\frac{n(n+1)}{2(2^{n+1}-2)}$ $<(1+2^{-20})\frac{1}{8}\sum_{n=21}^{\infty}\frac{n(n+1)}{2^{n-1}}$
$=(1+2^{-20})\frac{1}{8}\left(\frac{2}{(1-\frac{1}{2})^{3}}-\sum_{k=0}^{20}\frac{k(k+1)}{2^{k-1}}\right)$
$\approx\numprint{0.000121117}$
The product of this with $\delta_{1}-\delta_{0}$ is less than $\numprint{0.0002}$. We still need to take into account the contribution to the upper bound from the series with the $H_{n+1}$’s and the $\epsilon_{n}$’s. We expect something about one thousand times smaller than what we computed in the previous proof for the remainders, and indeed, one obtains:
[…]
$(\log 2)^{-1}\sum_{n=21}^{\infty}(1+2^{-n+1})\epsilon_{n}$ $<(\log 2)^{-1}(1+2^{-20})\sum_{n=21}^{\infty}2^{-n-1}$
[…]
As $\delta_{2}\approx\numprint{-0.311608}$ is, as we see from Table 3, the maximum among the $\delta_{m}$’s, for $m\neq 1$ and $0\leq m\leq 20$, we conclude, with some margin, that
$m\geq 2\implies\delta_{m}<\numprint{-0.3114}\;.$
Similarly, regarding the lower bound, we can separate out $\delta_{0}=-(\log 2)^{-1}$, writing
$(2^{p+1}-2)^{-1}\sum_{j=1}^{p}\binom{p+1}{j}\delta_{p-j}\geq\frac{p+1}{(2^{p+1}-2)}\delta_{0}+\min\{\delta_{j},1\leq j<p\},$
where we used the fact that we know $\delta_{j}\leq 0$ for all $j$’s. Now working again with $p>10$, we obtain (as the other contributions involving the $H_{n+1}$’s and the $\epsilon_{n}$’s are positive):
$p>10\implies\delta_{p}\geq\delta_{0}\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}+\underbrace{\min\{\delta_{j},1\leq j\leq 10\}}_{{}=\delta_{3}}.$
Now
$\sum_{n=11}^{\infty}\frac{n+1}{(2^{n+1}-2)}$ $<(1+2^{-10})\sum_{n=11}^{\infty}\frac{n+1}{2^{n+1}}$
$=(1+2^{-10})\frac{1}{2}\left(\frac{1}{(1-\frac{1}{2})^{2}}-\sum_{k=0}^{10}(k+1)2^{-k}\right)$
$\approx\numprint{0.006354}.$
The product of this with $\delta_{0}$ is about $\numprint{-0.009167}>\numprint{-0.01}$ and we thus can conclude that for all $m\geq 1$, $\delta_{m}>\numprint{-0.01}+\delta_{3}>\numprint{-0.35}$. ∎
4. The asymptotic behavior of $(e_{m})$
We have seen in the proof of Theorem 1 that we can define $e_{m}$ as $(m+1)c_{m}^{\prime}(1)/\log 2$ where the meromorphic functions $c_{m}(s)$ verify the recurrence Equation (4). From [5] and [4], there holds $c_{m}(s)=\frac{2^{s}-2}{2^{s}}u_{m}(s)$, where the meromorphic function $u_{m}(s)$ is, for $\Re s>1$, the $m$th moment of a certain discrete (complex) measure on $[0,1]$:
$\Re s>1\implies u_{m}(s)=\sum_{l=0}^{\infty}\biggl(\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-ls}\,.$ (10)
In particular $u_{0}(s)=\sum_{l=0}^{\infty}2^{-l(s-1)}=2^{s}/(2^{s}-2)$. As the function $t\mapsto t^{m}$ is non-decreasing and convex, there holds
$0\leq\int_{0}^{1}t^{m}\mathrm{d}t-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\leq\frac{1}{2}2^{-l}.$ (11)
So, at first for $\Re s=1$:
$u_{m}(s)$ $=(m+1)^{-1}\sum_{l=0}^{\infty}2^{-l(s-1)}-\sum_{l=0}^{\infty}\biggl((m+1)^{-1}-2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)u_{m}(s)$ $=\frac{2^{s}}{2^{s}-2}-\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}$
$(m+1)c_{m}(s)$ $=1-\frac{2^{s}-2}{2^{s}}\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)2^{-l(s-1)}.$
Thanks to estimate (11), this gives the analytic continution to $\Re s>0$, and we can now compute the value of the derivative at $s=1$:
$(m+1)c_{m}^{\prime}(1)=-\log(2)\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr).$
From Equation (7) which arose in the course of the proof of Theorem 1, we thus get an “explicit” formula for the coefficients $e_{m}$:
Proposition 2.
There holds:
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-(m+1)2^{-l}\sum_{0\leq n<2^{l}}\bigl(\frac{n}{2^{l}}\bigr)^{m}\biggr)\,.$ (12)
In terms of Bernoulli polynomials and numbers:
$e_{m}=\sum_{l=0}^{\infty}\biggl(1-\frac{B_{m+1}(2^{l})-B_{m+1}(0)}{2^{(m+1)l}}\biggr)=-\sum_{k=1}^{m}\binom{m+1}{k}B_{k}\frac{2^{k}}{2^{k}-1}\,.$ (13)
Remark 4.
The Bernoulli polynomials and numbers used in (13) are defined by $\sum_{n=0}^{\infty}B_{n}(x)\frac{t^{n}}{n!}=\frac{t\mathsf{e}^{xt}}{\mathsf{e}^{t}-1}$, $B_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}B_{k}x^{n-k}$, $B_{0}=1$, $B_{1}=-\frac{1}{2}$, $B_{n}=B_{n}(0)$, $(m+1)\sum_{0\leq n<N}n^{m}=B_{m+1}(N)-B_{m+1}(0)$. Formula (13) has large contributions of both signs and does not look promising numerically. In the sequel we will only use (12).
Remark 5.
A variant of (13) is (see Equation (22) for the Euler polynomials):
$e_{m}$ $=\sum_{l=0}^{\infty}(l+1)\frac{B_{m+1}(2^{l+1})-2^{m+1}B_{m+1}(2^{l})+(2^{m+1}-1)B_{m+1}(0)}{2^{(l+1)(m+1)}}$
$=\frac{m+1}{2}\sum_{l=0}^{\infty}(l+1)\frac{E_{m}(2^{l+1})-E_{m}(0)}{2^{(l+1)(m+1)}}\,.$
One only needs to write the numerator as $(l+1)(v_{l+1}-2^{m+1}v_{l})$ with $v_{l}=B_{m+1}(2^{l})-B_{m+1}(0)-2^{(m+1)l}$ and rearrange the sum. Details are left to the reader.
We will derive in this section from Equation (12) some asymptotic for $e_{m}$ as $m\to\infty$. From (11), the $l$th term is non-negative and bounded above by $(m+1)2^{-l-1}$. Let $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$. The sum of the contributions for $l>L_{m}$ is bounded above by $(m+1)2^{-L_{m}-1}$, hence, by $(m+1)^{-1}$.
To approximate the ($l$ dependent) moments $p_{l}=2^{-l}\sum_{0\leq n<2^{l}}(2^{-l}n)^{m}$ for $l\leq L_{m}$, we shall apply the same technique as in [6, Prop. 1]. For the convenience of the reader, the presentation here is entirely self-contained. [^1] In v1 of [6], which is the current version at time of writing, the Proposition 1 is formulated in a way which does not make explicit the implied constants in the big-O, although this is quite easy to extract from the proof there. As here we need to apply this to mutliple probability measures, we had anyhow to reproduce parts of the argument from [6] in order to control the uniformity. Certainly, the published version of [6] will make more explicit the big-O bound. Let us define
$q_{l}=2^{-l}\sum_{0\leq n<2^{l}}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\,.$ (14)
There holds $0\leq x\mathsf{e}^{-x}\leq\mathsf{e}^{-1}$ for $x\geq 0$, hence $x^{m}\leq\mathsf{e}^{-m(1-x)}$. In particular $(2^{-l}n)^{m}\leq\mathsf{e}^{-m(1-2^{-l}n)}$ and $0\leq p_{l}\leq q_{l}$. We split $q_{l}-p_{l}$ into two parts:
$q_{l}-p_{l}=2^{-l}(\sum_{0\leq n<2^{l-1}}+\sum_{2^{l-1}\leq n<2^{l}})\biggl(\mathsf{e}^{-m(1-2^{-l}n)}-(2^{-l}n)^{m}\biggr)$ (15)
We bound trivially the first part by
$2^{-l}\sum_{0\leq n<2^{l-1}}\mathsf{e}^{-m(1-2^{-l}n)}\leq\int_{0}^{\frac{1}{2}}\mathsf{e}^{m(x-1)}\mathrm{d}x=\frac{1}{m}\mathsf{e}^{-m/2}(1-\mathsf{e}^{-m/2}).$
This, and everything following applies also to non-integer positive $m$’s, the quantity $e_{m}$ can be defined for them using (12).
Regarding the second part in Equation (15), we use (as in the proof of [6, Prop. 1]) the bound
$m\geq 0,\;\frac{1}{2}\leq x\leq 1\implies\mathsf{e}^{-m(1-x)}-x^{m}\leq m(1-x)^{2}\mathsf{e}^{-m(1-x)}.$
So, for positive $m$:
$0\leq q_{l}-p_{l}\leq\frac{\mathsf{e}^{-m/2}}{m}+2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}$
(Note that if $l=0$, the sum above is empty and has value 0; so we now assume $l\geq 1$ and the bound we obtain also applies with $l=0$). Let $\psi_{m}(x)=\frac{\mathrm{d}}{\mathrm{d}x}\left((1-x)^{2}\mathsf{e}^{-m(1-x)}\right)=\bigl(m(1-x)-2\bigr)(1-x)\mathsf{e}^{-m(1-x)}$. We write:
$2^{-l}\sum_{2^{l-1}\leq n<2^{l}}m(1-2^{-l}n)^{2}\mathsf{e}^{-m(1-2^{-l}n)}=\frac{m}{8}\mathsf{e}^{-m/2}+m2^{-l}\sum_{2^{l-1}\leq n<2^{l}}\int_{\frac{1}{2}}^{2^{-l}n}\psi_{m}(y)\mathrm{d}y$
Using the probability measure $\mu_{l}=2^{-l}\sum_{0\leq n<2^{l}}\delta_{2^{-l}n}$, which has the property $\mu_{l}([y,1])\leq 1-y$ for every $y\in[0,1]$, we can rewrite the second summand on the right-hand side as:
$)\displaystyle m\int_{[\frac{1}{2},1\rparen}\int_{\frac{1}{2}}^{x}\psi_{m}(y)\mathrm{d}y\,\mu_{l}(\mathrm{d}x$) $\leq m\int_{\frac{1}{2}\leq y\leq x<1}\max(\psi_{m}(y),0)\mathrm{d}y\,\mu_{l}(\mathrm{d}x)$
$\leq m\int_{\frac{1}{2}}^{1}\max(\psi_{m}(y),0)(1-y)\mathrm{d}y$
$=m\int_{\frac{1}{2}}^{\max(1-\frac{2}{m},\frac{1}{2})}\bigl(m(1-y)-2\bigr)(1-y)^{2}\mathsf{e}^{-m(1-y)}\mathrm{d}y$
$=m^{-2}\int_{\min(2,m/2)}^{m/2}(t-2)t^{2}\mathsf{e}^{-t}\mathrm{d}t\leq 18\mathsf{e}^{-2}m^{-2}.$
We have obtained, for positive $m$ (with some loss of precision near zero, but we are only interested into $m\to\infty$):
$0\leq q_{l}-p_{l}\leq(m^{-1}+\frac{m}{8})\mathsf{e}^{-m/2}+\frac{18}{\mathsf{e}^{2}m^{2}}\leq C_{1}m^{-2},$ (16)
with $C_{1}=\max_{t\geq 0}(t+t^{3}/8)\mathsf{e}^{-t/2}+18\mathsf{e}^{-2}$. It turns out that the maximum is attained at $t\approx\numprint{5.04}$ and its value is $\approx\numprint{1.693}$ and $18\mathsf{e}^{-2}\approx\numprint{2.436}$ and $C_{1}<\numprint{4.13}$. Of course when $m$ is large $18\mathsf{e}^{-2}+\epsilon$ gives a tighter bound.
The proof applies to any positive $m$, and for $m$ a positive integer it becomes a statement involving the values of Bernoulli polynomials at $N=2^{l}$. That $N$ was a power of $2$ is not important. We formulate here a dedicated statement, as we could not find such property in the standard references.
Proposition 3.
For any two positive integers $m$ and $n$, the following uniform estimate holds:
$0\leq\frac{1-\mathsf{e}^{-m}}{n\bigl(\mathsf{e}^{mn^{-1}}-1\bigr)}-\frac{B_{m+1}(n)-B_{m+1}(0)}{(m+1)n^{m+1}}\leq\frac{5}{m^{2}}.$ (17)
Proof.
Let’s use $N$ rather than $n$ which has served earlier as summation index. Recall $\sum_{0\leq n<N}n^{m}=(m+1)^{-1}\bigl(B_{m+1}(N)-B_{m+1}(0)\bigr)$. The proof leading to (16) works identically with an arbitrary positive integer $N$ in the place of $2^{l}$, if $N$ is even. If $N\geq 3$ is odd, $N=2M+1$, $M\geq 1$, we have to be careful about how we split (the analog of) $q_{l}-p_{l}$ into two parts. Formerly we compared the first part with an integral from 0 to $\frac{1}{2}$ but now, using $0\leq n<\frac{1}{2}N$, i.e., $0\leq n\leq M$, we see that the rectangle basis extends a bit further to $\frac{1}{2}+\frac{1}{2N}$. We could improve that (for $m$ large) by bounding the contribution of $x=M/N$ pointwise by $\mathsf{e}^{-m/2}$ (as done in [6, Prop. 1] which handles general positive measures). But let us continue with $\int_{0}^{\frac{1}{2}+\frac{1}{2N}}\mathsf{e}^{m(x-1)}\mathrm{d}x<\mathsf{e}^{m/2N}m^{-1}\mathsf{e}^{-m/2}\leq\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}$, for the first part. For the second part, i.e. for the contributions of the points $x=N^{-1}n$, $M+1\leq n<N$, we can still use $x=\frac{1}{2}$ as reference point, there is no change, except that we could gain a little in the $\frac{m}{8}\mathsf{e}^{-m/2}$ term (as the weight of the half-interval is now not $1/2$, but $M/N$). Anyway, we end up with a slightly deteroriated final global upper bound $\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}+\frac{m}{8}\mathsf{e}^{-m/2}+18\mathsf{e}^{-2}m^{-2}$, which is $\leq C_{2}m^{-2}$ with $C_{2}=\max_{t\geq 0}(t\mathsf{e}^{-t/3}+\frac{t^{3}}{8}\mathsf{e}^{-t/2})+18\mathsf{e}^{-2}$. The maximum is again attained near $t=5$, it is found to be $\numprint{2.27}$ and $C_{2}<4.67$. ∎
As a corollary if we let both $m$ and $n$ go to infinity such that the ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
$\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\frac{q}{\mathsf{e}^{q}-1}\,.$
This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but (17) makes it easy (and gives an $O(m^{-1})$ error estimate). Reference [11], which uses the saddle-point method in the complex domain, provides full asymptotics for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
We now resume the study of the quantities $e_{m}$. Recall that we have set $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$ and already estimated
$0\leq\sum_{l>L_{m}}\bigl(1-(m+1)p_{l}\bigr)\leq\frac{1}{m+1},$
so, combining this with Equations (12) and (16):
$\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)\leq e_{m}\leq\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)+(L_{m}+1)(m+1)\frac{C_{1}}{m^{2}}+\frac{1}{m+1}\,.$
Let us recall Equation (14) defining $q_{l}$ and record our current result:
$e_{m}=\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\bigr)+O(\frac{\log m}{m}).$ (18)
This motivates the definition of auxiliary coefficients $f_{m}$ and $g_{m}$, for positive real $m$:
$f_{m}$ $=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)$ (19)
$g_{m}$ $=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}$ (20)
Clearly the series defining $f_{m}$ converges uniformly for $m$ bounded. We bound trivially the terms in the series defining $g_{m}$::
$x>0\implies\frac{x}{\mathsf{e}^{x}-1}=\frac{x\mathsf{e}^{-x/2}}{2\sinh(x/2)}\leq\mathsf{e}^{-x/2}$
$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$
[…]
$\phi(t)=-t-\frac{1}{2}+\sum_{l=0}^{\infty}\bigl(1-\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\bigr)-\sum_{l=-1}^{-\infty}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\,.$ (21)
The reason for the extra term $\frac{1}{2}$ is that this makes the average of $\phi$ vanish, as will be seen later. As shown previously the two series are uniformly convergent for $0\leq t\leq 1$, so $\phi$ is a continuous function. Here is a stronger (elementary) statement:
Proposition 4.
Equation (21) defines $\phi(t)$ as an analytic function of $t$ in the horizontal strip $|\Im t|<\frac{\pi}{2\log(2)}$.
Proof.
Details are left to reader: $\sum_{l=0}^{\infty}(1-\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1})$ defines a meromorphic function in the entire complex plane, with poles at the elements of $2\pi i\mathbb{Z}\setminus\{0\}$, so in particular if we substitute $z=2^{t}$, we obtain an analytic function in the strip $|\Im t|<\frac{\pi}{2\log 2}$. Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2}-\epsilon$, $0<\epsilon<\frac{\pi}{2}$, from bounding below the denominator and above the numerator. This proves that $\phi(t)$ is as stated. ∎
Returning to Equation (18), we get:
$e_{m}$ $=L_{m}+1-\frac{m+1}{m}(1-\mathsf{e}^{-m})\biggl(L_{m}+1-\sum_{0\leq l\leq L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\biggr)+O(\frac{\log m}{m})$
$=(1+O(\frac{1}{m}))\Bigl(f_{m}-\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\Bigr)+O(\frac{\log m}{m}).$
For $l>L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$, $2^{-l}m\leq m/(m+1)^{2}$, hence:
$\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)=O(\sum_{l>L_{m}}\frac{m^{2}}{4^{l}})=O(\frac{m^{2}}{4^{L_{m}}})=O(m^{-2}).$
This gives, for $m\to+\infty$:
$e_{m}=(1+O(\frac{1}{m}))\Bigl(\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-2})+O(m\mathsf{e}^{-2m})\Bigr)+O(\frac{\log m}{m}).$
As $\phi$ is periodic, it is bounded and we have obtained the proof of the main estimate as given in Theorem 2:
Proposition 5.
The quantities $(e_{m})$ defined by the recurrence (1), or, more generally, defined for all real positive $m$ by Equation (12), verify the asymptotic estimate, as $m\to+\infty$:
$e_{m}=\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(\frac{\log m}{m}),$
where $\phi(t)$ is the function defined by Equation (21). It is 11-periodic and analytic for $|\Im t|<\frac{\pi}{2\log(2)}$.
We prove in the next section that $\phi(t)$ has zero mean.
5. A Fourier series
We compute the Fourier series of the $1$-periodic function $\phi(t)$ from Equation (21). Let $L$ be some positive integer. Let $n\in\mathbb{Z}$. We start with
$\int_{0}^{1}\mathsf{e}^{-2\pi int}\sum_{l=-\infty}^{L}\frac{2^{t-l}}{\mathsf{e}^{2^{t-l}}-1}\mathrm{d}t=\int_{0}^{1}\mathsf{e}^{-2\pi int}\sum_{l=-L}^{\infty}\frac{2^{t+l}}{\mathsf{e}^{2^{t+l}}-1}\mathrm{d}t$
$=\int_{-L}^{+\infty}\mathsf{e}^{-2\pi int}\frac{2^{t}}{\mathsf{e}^{2^{t}}-1}\mathrm{d}t=\frac{1}{\log 2}\int_{2^{-L}}^{\infty}\mathsf{e}^{-2\pi in\log_{2}(x)}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}$
$=\frac{1}{\log 2}\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}+\frac{1}{\log 2}\int_{2^{-L}}^{1}x^{-2\pi in(\log 2)^{-1}}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x$
$+\frac{1}{\log 2}\begin{cases}(-2\pi in(\log 2)^{-1})^{-1}\biggl[x^{-2\pi in(\log 2)^{-1}}\biggr]_{2^{-L}}^{1}=0&(n\neq 0),\\ L\log(2)&(n=0).\end{cases}$
Let us first handle $n=0$. We thus have:
$c_{0}(\phi)\coloneq\int_{0}^{1}\phi(t)dt$ $=\lim_{L\to\infty}\biggl(-\frac{1}{2}-\frac{1}{2}+(L+1)-\frac{1}{\log 2}\int_{2^{-L}}^{\infty}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\biggr)$
$=\lim_{L\to\infty}\biggl(L-\frac{1}{\log 2}\bigl[\log(1-\mathsf{e}^{-x})\bigr]_{2^{-L}}^{\infty}\biggr)$
$=\lim_{L\to\infty}\log_{2}\bigl(2^{L}(1-\mathsf{e}^{-2^{-L}})\bigr)=0.$
Suppose now $n\neq 0$. We obtain
$c_{n}(\phi)$ $\coloneq\int_{0}^{1}\mathsf{e}^{-2\pi int}\phi(t)dt=\frac{1}{2\pi in}-\frac{1}{\log 2}\begin{aligned} \biggl(&\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\\ &+\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x\biggr)\end{aligned}$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\lim_{h\to 0^{+}}\begin{aligned} \biggl(&\int_{1}^{\infty}x^{-2\pi in(\log 2)^{-1}+h}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}\\ &+\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}+h}\Bigl(\frac{1}{\mathsf{e}^{x}-1}-\frac{1}{x}\Bigr)\mathrm{d}x\biggr)\end{aligned}$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\lim_{h\to 0^{+}}\biggl(\int_{0}^{\infty}x^{-2\pi in(\log 2)^{-1}+h}\frac{\mathrm{d}x}{\mathsf{e}^{x}-1}-\int_{0}^{1}x^{-2\pi in(\log 2)^{-1}+h-1}\mathrm{d}x\biggr)$
$=\frac{1}{2\pi in}-\frac{1}{\log 2}\biggl(\Gamma\bigl(1-2\pi in(\log 2)^{-1}\bigr)\zeta\bigl(1-2\pi in(\log 2)^{-1}\bigr)-\frac{1}{-2\pi in(\log 2)^{-1}}\biggr)$
$=-(\log 2)^{-1}\Gamma\bigl(1-\frac{2\pi in}{\log 2}\bigr)\zeta\bigl(1-\frac{2\pi in}{\log 2}\bigr).$
We have used the basic Riemann formula $\Gamma(s)\zeta(s)=\int_{0}^{\infty}x^{s-1}(\mathsf{e}^{x}-1)^{-1}\mathrm{d}x$ for $\Re s>1$. Using the usual bound $|\zeta(1+it)|=O_{t\to+\infty}(\log t)$ ([12, Thm. 3.5]), and $|\Gamma(1+it)|\sim\sqrt{2\pi|t|}\exp(-\frac{\pi}{2}|t|)$ from Stirling formula in the complex domain (or more expediently from the reflection formula $\Gamma(s)\Gamma(1-s)=\pi/\sin(\pi s)$), we recover the analyticity of $\phi(t)$ in the strip $|\Im t|<\pi/\log(4)$ which was stated in Proposition 4.
This completes the proof of Theorem 2.
6. Exponential generating function
Let us point out that there is an alternative formula for $f_{m}$, $m>0$, which is more convenient numerically:
Proposition 6.
Let $f_{m}$ for real $m>0$ be defined by equation (19),i.e.
$f_{m}=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr).$
There holds:
$f_{m}=\sum_{l=1}^{\infty}\frac{l2^{-l}m}{\mathsf{e}^{2^{-l}m}+1}\,.$
The author did not find the next formula in standard references [1, 7, 8] although it is undoubtly well-known, and is related to the relation
$E_{n-1}(x)=2n^{-1}\bigl(B_{n}(x)-2^{n}B_{n}(2^{-1}x)\bigr)$ (22)
between Euler and Bernoulli polynomials ([1, 23.1.27], [7, 1.14(7)]).
Lemma 2.
For $z\in\mathbb{C}\setminus 2\pi i\mathbb{Z}$,
$\sum_{k=1}^{\infty}\frac{2^{-k}z}{\exp(2^{-k}z)+1}=1-\frac{z}{\mathsf{e}^{z}-1}\,.$ (23)
Observe in passing that if we multiply both sides by $\mathsf{e}^{z}-1$ we obtain an identity of entire functions.
Proof.
For any integer $l\geq 1$ we consider the identity
$\mathsf{e}^{z}-1=(e^{z/2}+1)(\mathsf{e}^{z/4}+1)\cdots(\mathsf{e}^{z/2^{l}}+1)(\mathsf{e}^{z/2^{l}}-1),$
and compute its logarithmic derivative. This gives, for $z$ not among the poles:
$\frac{\mathsf{e}^{z}}{\mathsf{e}^{z}-1}=\sum_{k=1}^{l}\frac{\mathsf{e}^{z/2^{k}}2^{-k}}{\mathsf{e}^{z/2^{k}}+1}+\frac{\mathsf{e}^{z/2^{l}}}{2^{l}(\mathsf{e}^{z/2^{l}}-1)}.$
The limit for $l\to\infty$ gives:
$1+(\mathsf{e}^{z}-1)^{-1}=\sum_{k=1}^{\infty}\bigl(2^{-k}-\frac{2^{-k}}{\mathsf{e}^{z/2^{k}}+1}\bigr)+z^{-1},$
which, after canceling out the $1$ and multiplying by $z$ gives Equation (23). ∎
Proof of Proposition 6.
We can now compute:
$f_{m}$ $=\sum_{l=0}^{\infty}\Bigl(1-\frac{m/2^{l}}{\exp(m/2^{l})-1}\Bigr)$
$=\sum_{l=0}^{\infty}\sum_{k=1}^{\infty}\frac{2^{-l-k}m}{\exp(2^{-l-k}m)+1}$
$=\sum_{q=1}^{\infty}(\sum_{k=1}^{q}1)\frac{2^{-q}m}{\exp(2^{-q}m)+1}\,.\qed$
The statement and proof of Proposition 6 actually hold for any complex $m$ which is not in $2\pi i\mathbb{Z}$.
In the next proposition we compute (the primitive of) the exponential generating function for the $(e_{m})_{m\geq 0}$ sequence.
Proposition 7.
There holds (for $t\in\mathbb{C}$):
$\sum_{m=0}e_{m}\frac{t^{m+1}}{(m+1)!}=(\mathsf{e}^{t}-1)\sum_{l=0}^{\infty}\Bigl(1-\frac{2^{-l}t}{\mathsf{e}^{2^{-l}t}-1}\Bigr)=(\mathsf{e}^{t}-1)\sum_{l=1}^{\infty}\frac{l2^{-l}t}{\mathsf{e}^{2^{-l}t}+1}\,.$ (24)
Proof.
Note that the quantity $E(t)$ defined by the left-hand side is an entire function, and indeed if we move the factor $\mathsf{e}^{t}-1$ to inside the sums we see that the two other expressions are, too. Let us use as starting point Equation (13) (and recall $e_{0}=0$):
$\sum_{m=1}e_{m}\frac{t^{m+1}}{(m+1)!}$ $=-\sum_{m=1}\frac{t^{m+1}}{(m+1)!}\sum_{k=1}^{m}\frac{(m+1)!}{k!(m+1-k)!}B_{k}\frac{2^{k}}{2^{k}-1}$
$=-\sum_{k=1}^{\infty}\frac{B_{k}2^{k}}{k!(2^{k}-1)}\sum_{m=k}^{\infty}\frac{t^{m+1}}{(m+1-k)!}$
$=-\sum_{k=1}^{\infty}\frac{B_{k}2^{k}t^{k}}{k!(2^{k}-1)}(\mathsf{e}^{t}-1)$
$=-(\mathsf{e}^{t}-1)\sum_{k=1}^{\infty}B_{k}\frac{t^{k}}{k!}\sum_{l=0}^{\infty}2^{-lp}$
$=-(\mathsf{e}^{t}-1)\sum_{l=0}^{\infty}\Bigl(\frac{2^{-l}t}{\mathsf{e}^{2^{-l}t}-1}-1\Bigr).$
This gives the first equality from Equation (24). The second one was already stated, in another variable, as Proposition 19. ∎
Remark 6.
Hence, if we let $E(t)$ denote the left-hand side of Equation (24), we see that $E(m)=(\mathsf{e}^{m}-1)f_{m}=(\mathsf{e}^{m}-1)\bigl(e_{m}+O(\frac{\log m}{m})\bigr)=\mathsf{e}^{m}e_{m}\bigl(1+O(m^{-1})\bigr)$ and thus $e_{m}=(1+O(m^{-1}))\mathsf{e}^{-m}E(m)=\mathsf{e}^{-m}E(m)+O(\frac{\log m}{m})$. This is very reminiscent of the general estimate from [6, Prop. 1] relating moments of some measures on the unit interval with their exponential generating function.
Remark 7.
A form of the functional equation verified by $E(t)$ (which has neither constant term nor a linear term, as $e_{0}=0$) is
$E(2t)=(\mathsf{e}^{t}+1){E(t)}+\mathsf{e}^{2t}-1-2t.$ (25)
Looking at (25) as an identity of formal power series, it implies that $E$ has no constant term, but not that $E^{\prime}(0)=0$. Equation (25) with the definition $E(t)=\sum_{m=0}^{\infty}e_{m}\frac{t^{m+1}}{(m+1)!}$ is equivalent to the recurrence relation Equation (1) but stated without the initial condition $e_{0}=0$.
7. Concluding remarks
The $\phi$ function from Theorem 2 fits into an old theme going back to Hardy [9] (see also [10]). Balazard et al. [3] have studied under some general conditions series of the type $f_{\alpha}(x)=\sum_{n\in\mathbb{Z}}\alpha(x^{\theta^{n}})$ where $\alpha$ is continuous on $[0,1]$, vanishes at 0 and $1$ and has a power series expansion $\alpha(x)=\sum_{k=1}^{\infty}a_{k}x^{k}$ (convergent for $0\leq x<1$). They obtain the Fourier series of $f_{\alpha}$ as a function of $\log_{\theta}\log x^{-1}$ in terms of values at $(\log\theta)^{-1}2\pi i\mathbb{Z}\setminus\{0\}$ of the Gamma function and a Dirichlet series associated to $\alpha$. This is reminiscent of the computations done in section 5 (we did not need to use the Poisson summation formula).
They consider in particular the case $\alpha(x)=x\prod_{n=0}^{\infty}(1-x^{2^{n}})$, which is related to the Thue-Morse sequence (see also about this [2]). It is of note that the function $\phi(t)$ ($t=\log_{2}(m)$) from Theorem 2 and Equation (21) is related to the logarithmic derivative of $m\mapsto\mathsf{e}^{m}\alpha(\mathsf{e}^{-m})$. Indeed, we compute straightforwardly:
$m\frac{\mathrm{d}}{\mathrm{d}m}\log\prod_{n=0}^{\infty}(1-\mathsf{e}^{-2^{n}m})=m\sum_{n=0}^{\infty}\frac{2^{n}\mathsf{e}^{-2^{n}m}}{1-e^{-2^{n}m}}=\sum_{n=0}^{\infty}\frac{2^{n}m}{\mathsf{e}^{2^{n}m}-1}=\frac{m}{\mathsf{e}^{m}-1}-g_{m}\,.$
Here $g_{m}$ is the quantity defined in Equation (20).
This computation motivates the consideration of the following infinite product:
$\Psi(z)=2^{\frac{\operatorname{Log}_{2}(z)(\operatorname{Log}_{2}(z)+1)}{2}}\prod_{l=0}^{\infty}\frac{1-\mathsf{e}^{-2^{-l}z}}{2^{-l}z}\prod_{k=1}^{\infty}(1-\mathsf{e}^{-2^{k}z}).$ (26)
The first infinite product is an entire function. The second one is analytic in the half-plane $\Re z>0$. So $\Psi$ is analytic there. It verifies the functional equation:
$\Psi(2z)=2^{\frac{(\operatorname{Log}_{2}(z)+1)(\operatorname{Log}_{2}(z)+2)}{2}-\frac{\operatorname{Log}_{2}(z)(\operatorname{Log}_{2}(z)+1)}{2}}\Psi(z)\frac{(1-\mathsf{e}^{-2z})/(2z)}{1-\mathsf{e}^{-2z}}=\Psi(z)$
A straightforward computation gives the logarithmic derivative with respect to the variable $\operatorname{Log}z$:
$z\frac{\mathrm{d}}{\mathrm{d}z}\log\Psi(z)=\operatorname{Log}_{2}(z)+\frac{1}{2}+\sum_{l=0}^{\infty}\Bigl(\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1}-1\Bigr)+\sum_{k=1}^{\infty}\frac{2^{k}z}{\mathsf{e}^{2^{k}z}-1}.$
Comparison with Equation (21) shows that the above is exactly $\phi(\operatorname{Log}_{2}z)$. And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an nalytic function.
Whether this relation between $\Psi$ and $\phi$ is indicative of some closer connection between the sequence $(e_{m})$ and the Thue-Morse sequence [2, 3] is currently unknown to the author.